nt Assume the firm finds that its profit will be at maximum when it produces $400 worth of product A. Suppose also that each of the two techniques shown in the following table will produce the desired output. If the company choose the most desirable technique, will the industry expand or contract? Price Resources Resource units required resources Technique Technique Technique Technique Labor Land Capital Entrepreneurial $2 $4 $3 $5 #1 50 20 20 40 #2 30 20 50 40 #3 40 30 30 50 #4 35 20 40 45 Expand O Contract O both either
Answer #1

Resource unit required

Cost of technique

Resources

Price resources

Technique 1

Technique 2

Technique 3

Technique 4

Technique 1

Technique 2

Technique 3

Technique 4

Labor

2

50

30

40

35

100

60

80

70

Land

4

20

20

30

20

80

80

120

80

Capital

3

20

50

30

40

60

150

90

120

Entrepreneurial

5

40

40

50

45

200

200

250

225

440

490

540

495

Calculation:

Cost of technique

Technique 1

Technique 2

Technique 3

Technique 4

2*50=100

2*30=60

2*40=80

2*35=70

4*20=80

4*20=80

4*30=120

4*20=80

3*20=60

3*50=150

3*30=90

3*40=120

5*40=200

5*40=200

5*50=250

5*45=225

100+80+60+200=440

60+80+150+200=490

80+120+90+250=540

70+80+120+225=495

Technique 1

440

Technique 2

490

Technique 3

540

Technique 4

495

The firm will choose Technique 1 because it produces the output at the least cost. (Compared to other techniques) Production using technique 1 will give loss, Since TC > TR

TR = $400

TC = $440

Loss = TC – TR = 440-400

Loss = $40

Loss will be $40, causing the industry to contract.

Answer: Contract